# ControlMM: Controllable Masked Motion Generation

### v1

@@ 3.3 Inference-time Logits and Codebook Editing
At each iteration $i$, the logits $l_{i}$ are updated using the following gradient-based approach:
$l_{i+1}=l_{i}-\eta\nabla_{l_{i}}L_{s}(l_{i},s).$ (7)
where $\eta$ controls the magnitude of the updates to the logits, while $L_{s}(l_{i},s)$ represents the gradient of the objective function with respect to the logits $l$ at iteration $p$. This refinement process continues over $P$ iterations. Similarly, we can also edit the entries in motion codebook to minimize the consistency loss:
$e_{c}^{i+1}=e_{c}^{i}-\eta\nabla_{e_{c}^{i}}L_{s}(e_{c}^{i},s),$ (8)
where $e_{c}$ represents the embedding in the codebook space. Our experiments demonstrate that combining joint logits and codebook editing results in the best performance.
@@ 4 Applications
Any Joints Any Frame Control . To control specific joints at particular frames, the spatial control signal can be directly applied to the desired joint and frame in the global position, as the loss function during training is specifically designed for this task.
Obstacle Avoidance. Since inference-time logits and codebook editing is versatile, it can be compatible with arbitrary loss function. The Signed Distance Function (SDF) can serve as a loss function for obstacle avoidance, where the gradient field dictates the direction to repel from obstacles. This loss function incorporates a safe distance threshold $d$, beyond which the gradient diminishes to zero, and is defined as:
$\mathcal{L}_{\text{obs}}(x):=\sum_{i,n}-\min\left[\text{SDF}\left(\hat{c}_{i,n}(x)\right),d\right],$ (9)
where $\text{SDF}_{n}$ denotes the Signed Distance Function for obstacle $i$ in frame $n$, which can change across frames in the case of moving obstacles. While this application is similar to the one proposed by GMD (Karunratanakul et al., 2023), ControlMM offers enhanced functionality by enabling obstacle avoidance for any joint at any frame, rather than being limited to the root trajectory (pelvis) as proposed in GMD.
Body Part Timeline Control. ControlMM supports motion generation conditioned on multiple joints, enabling control over body parts. To support multiple prompts corresponding to various body parts and timelines, ControlMM processes each prompt sequentially. Initially, it generates motion without any body part control, then iteratively refines the motion by incorporating prompts conditioned on the specified body parts and timeline constraints from the prior generation. Since ControlMM allows spatial control signals to target any joint and frame, partial body or temporal frame control is applicable within this framework. The detail of this process is described in A.9.
@@ 5 Experiment
Datasets. We conduct comprehensive experiments on the HumanML3D dataset (Guo et al., 2022b) HumanML3D covers a wide variety of everyday human actions, such as exercising and dancing. It includes 14,616 motion sequences accompanied by 44,970 text descriptions. The textual data contains 5,371 unique words. The motion sequences are sourced from AMASS (Mahmood et al., 2019) and HumanAct12 (Guo et al., 2020), with each sequence downsampled to 20 frames per second (FPS) and trimmed to a duration of 10 seconds, yielding clips between 2 and 10 seconds long.
Evaluation. We follow the evaluation protocol from OmniControl (Xie et al., 2023) which combines evaluation of quality from HumanML3D(Guo et al., 2022b) and trajectory error from GMD (Karunratanakul et al., 2023). The Frechet Inception Distance (FID) is used to assess the naturalness of the generated motion. R-Precision measures how well the generated motion aligns with its corresponding text prompt, while Diversity captures the variability within the generated motion. To assess control performance, we use the foot skating ratio, following Karunratanakul et al. (2023), as an indicator of coherence between the motion trajectory and the physical plausibility of the human motion. We also report Trajectory error, Location error, and Average error of the controlled joint positions in keyframes to evaluate control accuracy. All models are trained to generate 196 frames for evaluation, using 5 levels of sparsity in the control signal: 1, 2, 5, 49 (25% density), and 196 keyframes (100% density). Keyframes are sampled randomly, and we report the average performance across all density levels. During both training and evaluation, models receive ground-truth trajectories as spatial control signals.
5.1 Quantitative Comparison to State-of-the-art Approaches
@@
Table 1: Comparison of text-condition motion generation with spatical control signal on the HumanML3D. The first section, “Train on Pelvis Only,” evaluates our model that was trained solely on the pelvis. The last section, “Train on All Joints”, is trained on all joints and assessing performance for each one. The cross-section reports performance across various combinations of joints.

### v2

@@ 3.3 Logits Optimization
At each iteration $m$, the logits $l_{m}$ are updated using the following gradient-based approach:
$l_{m+1}=l_{m}-\eta\nabla_{l_{m}}L_{s}(l_{m},s),$ (7)
where $\eta$ controls the magnitude of the updates to the logits, while $L_{s}(l_{m},s)$ represents the gradient of the objective function w.r.t. the logits $l_{m}$. This refinement process continues over $I$ iterations. Similarly, in the last unmask step, optimizing embeddings from the codebook space directly is possible. We can directly optimize the embedding:
$e_{m+1}=e_{m}-\eta\nabla_{e_{m}}L_{s}(e_{m},s),$ (8)
where $e$ represents the embedding in the codebook space. For zero-shot objective control, loss function $L_{s}$ can be replaced with any arbitrary objective function.
3.4 Differentiable Expectation Sampling (DES)
Differentiable Sampling. Both logits regularizer and optimization update the logits by computing the gradient of the disparity between spatial-temporal control signals and generated motion sequence defined in eq. 3. This requires sampling motion tokens according to the categorical token distribution during training/optimization time, which is inherently non-differentiable. To address this, we apply Straight-Through Gumbel-Softmax Estimator [19], a reparameterization trick that allows differentiable sampling from a categorical distribution, i.e.,
$p_{\theta}\left(x_{k}\mid X_{\overline{\mathbf{M}}},W,S\right)=\frac{\exp\left((\ell_{k}+g_{k})/\tau\right)}{\sum_{j=1}^{K}\exp\left(\ell_{j}+g_{j}/\tau\right)},$ (9)
where $l$ is logits, $\tau$ refers to temperature, and $g$ represents Gumbel noise with $g_{1},\dots,g_{K}$ being independent and identically distributed (i.i.d.) samples from a $\text{Gumbel}(0,1)$ distribution. The $\text{Gumbel}(0,1)$ distribution can be sampled via inverse transform sampling by first drawing $u\sim\text{Uniform}(0,1)$ and then computing $g=-\log(-\log(u))$.
Token Expectation. Masked models reconstruct an embedding from the reconstruction probability $p_{\theta}(x_{k}\mid\cdot)$ by querying the discrete token embedding using the index with the highest probability, $\operatorname{argmax}_{k}p_{\theta}(x_{k}\mid\cdot)$. However, $\operatorname{argmax}$ is non-differentiable, and cannot be controlled by perturbation from a control signal. To address this, we relax the quantization constraint by utilizing the weighted average of the codebooks (or transformer tokens) w.r.t. the reconstruction probability $p_{\theta}(x_{k}\mid\cdot)$ to approximate continuous tokens. This can be expressed as the expectation of the reconstructed embedding $X_{recon}$:
$\mathbb{E}[X_{recon}]=\sum_{k=1}^{K}p_{\theta}\left(x_{k}\mid X_{\overline{\mathbf{M}}},W,S\right)\cdot c_{k}.$ (10)
Figure 2: Visualization comparisons to state-of-the-art methods for any-joint any-frame control. The plots on the top display the top view of pelvis control (root trajectory), while the bottom plot shows the side view of the right wrist. Red represents the control signal, and Blue represents the generated joint motion.
Figure 3: Visualization comparisons to state-of-the-art methods for zero-shot objective control. Objective: constrain a human to walk inside a square area.
@@ 4 Applications
Any-Joints-Any-Frame Control. To control specific joints at particular frames, the joint control signal can be directly applied to the desired joint and frame in the global position, as the loss function during training is specifically designed for this task.
Body-Part Timeline Control. MaskControl supports motion generation conditioned on multiple joints, enabling control over body parts. To handle multiple prompts corresponding to various body parts and timelines, MaskControl processes each prompt sequentially. Initially, it generates motion without any body part control, then iteratively refines the motion by incorporating prompts conditioned on the specified body parts and timeline constraints from the prior generation. Since MaskControl allows joint control signals to target any joint and frame, partial body or temporal frame control is applicable within this framework. The detail of this process is described in A.12.
Zero-shot Objective Control. In many human motion generation tasks, motion generation requires adaptation to dynamic constraints at inference time, such as human-scene interaction, human-object interaction, or human self-contact. Logits Optimization allows optimization with arbitrary loss functions that take joints and frames as inputs, enabling zero-shot adaptation at inference time. However, direct optimization to satisfy arbitrary objective functions can often lead to unrealistic motion. To mitigate this, Logits Optimization perturbs the logits during the unmasking process, allowing the masked transformer to re-predict them. By doing so, it ensures that the generated motion remains close to the learned distribution while still achieving the desired control.
@@ 5 Experiment
Table 2: Comparison of text-condition motion generation with joint control signal on the HumanML3D. The first section, “Train on Pelvis Only,” evaluates our model that was trained solely on the pelvis. The last section, “Train on All Joints”, is trained on all joints and reports the average evaluation for each joint. $\rightarrow$ indicates the closer to the real value, the better.
Method Joint R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Trajectory Error (>50 cm) () $\downarrow$ Location Error (>50 cm) () $\downarrow$ Average Error (cm) $\downarrow$
Real - 0.797 0.002 9.503 - 0.00 0.00 0.00
[…]
MDM 0.602 0.698 9.197 0.1019 40.22 30.76 59.59
PriorMDM 0.583 0.475 9.156 0.0897 34.57 21.32 44.17
GMD 0.665 0.576 9.206 0.1009 9.31 3.21 14.39
OmniControl (on pelvis) Pelvis 0.687 0.218 9.422 0.0547 3.87 0.96 3.38
MotionLCM 0.752 0.531 9.253 - 18.87 7.69 18.97
TLControl 0.779 0.271 9.569 - 0.00 0.00 1.08
Ours (on pelvis) 0.809 0.061 (-77%) 9.496 0.0547 0.00 0.00 0.98
[…]
OmniControl  0.693 0.310 9.502 0.0608 6.17 1.07 4.04
TLControl Average 0.782 0.256 9.719 - 0.00 0.00 1.11
Ours 0.805 0.083 (-68%) 9.395 0.0545 0.00 0.00 0.72
Datasets. We conduct comprehensive experiments on the HumanML3D dataset [15] HumanML3D covers a wide variety motions. It includes 14,616 motion sequences accompanied by 44,970 text descriptions. The textual data contains 5,371 unique words. The motion sequences are sourced from AMASS [34] and HumanAct12 [13].
Evaluation. We follow the evaluation protocol from OmniControl [55] which combines evaluation of quality from HumanML3D[15] and trajectory error from GMD [21].
5.1 Quantitative Comparison to State-of-the-art
GMD [21] only addresses the pelvis location on the ground plane (xz coordinates). To ensure a fair comparison, we follow OmniControl [55] and compare GMD in managing the full 3D position of the pelvis (xyz coordinates). The first section of Tab. 2 resents results for models trained on the pelvis alone to ensure a fair comparison with previous state-of-the-art methods on the HumanML3D [15] dataset. $\rightarrow$ means closer to real data is better. Our model demonstrates significant improvements across all evaluation metrics. When compare to TLControl, the FID score notably decreased from 0.271 to 0.061, the R-Precision increased from 0.779 to 0.809, indicating superior generation quality. In terms of joint control accuracy, both Trajectory Error and Location Error dropped to zero, while the average error decreased to 0.91 cm, indicating highly precise joint control. Furthermore, our model outperforms existing methods in both Diversity and Foot Skating Ratio metrics. In the second section, Train on All Joints, we follow the evaluation from OmniControl [55], as our model supports control of any joint, not just the root (pelvis). We train the model to control multiple joints, specifically the pelvis, left foot, right foot, head, left wrist, and right wrist. Our model constantly outperform SOTA in all metrics.
5.2 Qualitative Comparison to State-of-the-art
Any-joint-Any-Frame. We visualize the generated motion using GMD [21], MotionLCM [7] and OmniControl [55] in Fig. 2. The motion is generated based on the prompt “a person walks forward and waves his hands,” with the pelvis and right wrist controlled in a zigzag pattern. Since GMD can only control the pelvis, we apply control only to the pelvis for GMD. However, it fails to follow the zigzag pattern, tending instead to move in a straight line. OmniControl receives control signals for both the pelvis and right wrist. Yet, it not only fails to follow the root trajectory (pelvis) but also does not adhere to the zigzag pattern for the right wrist. In contrast, our MaskControl demonstrates realistic motion with precise joint control for both the pelvis and the right wrist, accurately following the intended zigzag pattern.
Zero-shot Objective Control. Fig. 3 shows the visualization results with the unseen objective function, ‘Constrain a human to walk inside a square area,’ compared to ProgMoGen [30]. ProgMoGen not only violates the constraint but also generates motion unrelated to the text, as it lacks any ‘turns around’ motion.
5.3 Body Part Editing
Top1 Top2 Top3
MDM [48] 0.298 0.462 0.571 4.827 4.598 7.010
OmniControl [55] 0.374 0.550 0.656 1.213 5.228 9.258
MMM [40] 0.500 0.694 0.798 0.103 2.972 9.254
MotionLCM [7] 0.512 0.685 0.798 0.311 2.948 9.736
Ours 0.517 0.708 0.804 0.074 2.945 9.380
With joint signal controls, our model is capable of conditioning on multiple joints, which can be treated as distinct body parts, while generating the remaining body parts based on text input. In Tab. 3 We quantitatively compare our approach to existing methods designed for this task, including MDM [48] and MMM [40]. Additionally, we compare it with OmniControl [55] and MotionLCM [7], which also supports joint signal control. However, our evaluation demonstrates that OmniControl performs poorly in this task. Following the evaluation protocol from [40], we condition the lower body parts on ground truth for all frames and generate the upper body based on text descriptions using the HumanML3D dataset [15]. Our model is evaluated without retraining, using the same model as in the Train on All Joints setup, ensuring a fair comparison with OmniControl, which is trained on a subset of joints. Specifically, we condition only on the pelvis, left foot, and right foot as the lower body signals.
The results show that MDM struggles significantly when conditioned on multiple joints, with the FID score increasing to 4.827. Although OmniControl supports multiple joint control, our experiments reveal that it also suffers under these conditions, with its FID score rising to 1.213. This is consistent with the Cross-Joint evaluation in Tab. 7, which evaluate on multiple joint combination, where OmniControl’s FID score deteriorates considerably. MMM performs well in this task but requires retraining with separate codebooks for upper and lower body parts. In contrast, our model outperforms all other methods across all metrics without any retraining. When comparing to the ‘Train on Pelvis Only’ setup in Tab. 2, our model achieves similar FID and R-Precision scores, highlighting its robustness in handling multiple joint control signals.
Table 4: Comparison of zero-shot objective control. Three Human-Scene Interaction objectives are adopted from the programmable motion model (ProgMoGen [30]). Both ProgMoGen and MaskControl are able to control motion during inference by arbitrary loss functions, while MDM and MoMask serve as uncontrollable baseline models.
Task HSI-1: Head Height Constraint
Method Foot Skate $\downarrow$ Max Acc. $\downarrow$ Constraint Error $\downarrow$ Unsucc. Rate $\downarrow$ FID $\downarrow$ Diversity $\rightarrow$ R-prec. (Top3) $\uparrow$
[…]
ProgMoGen [30] 0.075 0.094 0.012 0.088 0.556 9.611 0.597
Ours 0.066 0.074 0.000 0.000 0.246 9.393 0.695
Task HSI-2: Avoiding Barrier Task HSI-3: Walking Inside a Square
Method Foot Skate $\downarrow$ Max Acceleration $\downarrow$ Constraint Error $\downarrow$ Foot Skate $\downarrow$ Max Acceleration $\downarrow$ Constraint Error $\downarrow$
[…]
ProgMoGen [30] 0.189 0.150 0.097 0.125 0.093 0.012
Ours 0.146 0.126 0.000 0.092 0.078 0.000
5.4 Zero-shot Objective Control
We evaluate three zero-shot objective control tasks using three Human-Scene Interaction evaluations from ProgMoGen [30], a programmable motion control model, i.e. ‘Head Height Constraint’, ‘Avoiding Barrier’, and ‘Walking Inside a Square’. MDM [48] and MoMask [16] serve as baselines without any control mechanisms. Since ProgMoGen builds on MDM as its base model. Similarly MoMask is shown as a baseline for the unconstrained Mask Motion Model. Tab. 4 shows that our MaskControl outperform ProgMoGen in all metrics. Max Acc. represents the maximum acceleration of joints which indicates the amount of joint jittering. Constraint Error measures how much the head height exceeds the specified constraint. Unsucc. Rate represents the rate of motion sequences that violate the height constraint.
6 Ablation Study
6.1 Component Analysis
Table 5: Ablation results of components analysis and different densities of joint control signal.
Method R-Prec. Top-3 $\uparrow$ FID $\downarrow$ Foot Skat. $\downarrow$ Traj. Err. $\downarrow$ Loc. Err. $\downarrow$ Avg. Err. $\downarrow$
[…]
No Control 0.807 0.095 0.0527 50.66 35.11 63.18
w/o Logits Regularizer 0.795 0.142 0.0577 0.32 0.02 2.18
w/o Logits Optimization 0.802 0.128 0.0594 39.14 24.00 40.41
Full model 0.809 0.061 0.0547 0.00 0.00 0.98
[…]
Density: 1 0.804 0.077 0.0551 0.00 0.00 0.10
Density: 2 0.806 0.087 0.0553 0.00 0.00 0.34
Density: 5 0.811 0.078 0.0553 0.00 0.00 0.98
Density: 25% 0.812 0.055 0.0536 0.01 0.00 1.68
Density: 100% 0.814 0.054 0.0543 0.02 0.00 1.64
To understand how each component impact the quality and joint control error. We conduct an ablation study on each component in sec. Component Analysis of Tab. 5, using same evaluation as Tab. 2. Without any control, No Control, the model achieves the highest diversity and the lowest Foot Skating Ratio, indicating strong realism in the generated motion. The FID score is also on par. However, all spatial errors are poor due to the absence of joint control components in the model. Without the Logits Regularizer, the average error remains relatively low, but the FID score is the worst, highlighting the importance of the Logits Regularizer for generation quality. Without Logits Optimization, the average error significantly worsens, although the FID remains acceptable, suggesting that the Logits Regularizer helps maintain generation quality. In the Full model, both the Logits Regularizer and Logits Optimization complement each other, improving both generation quality (FID) and control accuracy (Average Error).
6.2 Density of Joint Control Signal
In second part of table 5, we provide a detailed analysis of MaskControl’s performance across five different joint control density levels, where the model is trained for pelvis control using the HumanML3D dataset. The values 1, 2, and 5 refer to controlling the motion using exactly 1, 2, or 5 frames. The values 25% and 100% indicate the percentage of the total ground-truth motion length for each sample, which can range from 40 to 196 frames depending on the sample. The results show that increasing the joint control improves the quality: the FID score decreases from 0.077 with 1-frame control to 0.054 with full 196-frame (100%) control. Similarly, R-Precision improves from 0.804 at 1-frame density to 0.814 at 196-frame (100%) density. However, the Average Error shows the opposite trend—more joint control leads to higher error, as the model is required to target more specific points.
7 Conclusion
In this work, we present MaskControl, the first model to introduce controllability to the generative masked motion model, enabling precise control over quantized motion tokens while maintaining high-quality motion generation consistently outperforming diffusion-based controllable frameworks. MaskControl introduces Differentiable Expectation Sampling (DES) to relax codebook quantization constraint, enabling two key innovations: Logits Regularizer uses random masking and reconstruction to ensure that the generated motions are of high fidelity, while also reducing inconsistencies between the input control signals and the motions produced. Inference-Time Logit Optimization fine-tunes the predicted motion distribution during the unmasking process to better match the input control signals, enhancing precision and making MaskControl adaptable for unseen tasks. MaskControl has a wide range of applications, including any-joint-any-frame control, body-part timeline control, and zero-shot objective control.
Appendix A Appendix
A.1 Overview
  • Section A.2: Comparision to STMC
  • Section A.3: Pseudo Code of MaskControl Inference
  • Section A.4: Implementation Details
  • Section A.5: Full Evaluation on All Joint
  • Section A.6: Inference speed, quality, and errors Details
  • Section A.7: Speed of each component
  • Section A.8: Quantitative result for all joints of MaskControl-Fast
  • Section A.9: Ablation on less number of generation steps
  • Section A.10: Analysis of Logits Optimization and Logits Regularizer
  • Section A.11: Ambiguity of Motion Control Signal
  • Section A.12: Body Part Timeline Control
  • Section A.13: Cross Combination
Video visualization can be found at https://www.ekkasit.com/ControlMM-page/
A.2 Comparision to STMC
The STMC [38] setting involves different controllable joints and longer motion sequences, making direct comparison infeasible without re-training the Logits Regularizer. Fortunately, Logits Optimization enables zero-shot objective control, allowing us to achieve the STMC setting without retraining. We adopt the following steps: (1) Generate motions for each prompt: Motions are generated separately without joint control. (2) Resolving unassigned timeframes: Following STMC, we apply the SINC heuristic to fill in unassigned joints based on the sparse body-part timeline input. (3) Any-Joint-Any-Frame Control by each body part: We use the control conditions from steps (1) and (2), and remove joint control from connected timeframes to create padding, enabling smooth transitions between adjacent motion segments. Finally, the full motion is generated using these joint control signals without any text prompt, effectively generating body-part timelines onto a blank canvas.
Figure 4: Generating body parts timeline for STMC setting.
We evaluate two variants: ‘No Control Pad’, where no control is applied in the padding regions, and ‘Merge Control Pad’, where the optimization weight gradually decreases from adjacent body-part controls. Our method outperforms STMC across all metrics, as shown in Tab. 6. Note that performance could be further improved by re-training the Logits Regularizer for the STMC setting.
Per-crop semantic correctness Realism
Method R@1 $\uparrow$ R@3 $\uparrow$ TMR-Score $\uparrow$ FID $\downarrow$ Transition
M2T M2M distance $\downarrow$
Ground truth 55.0 73.3 0.748 1.000 0.000 1.5
STMC-MotionDiffuse 24.8 46.7 0.660 0.632 0.531 1.5
STMC-MDM 25.1 46.0 0.641 0.633 0.606 2.4
Our (No Control Pad) 38.3 58.1 0.688 0.654 0.511 1.6
Our (Merge Control Pad) 34.8 55.6 0.675 0.653 0.508 1.5
Table 6: Quantitative comparison with STMC
A.3 Pseudo Code of MaskControl Inference
Algorithm 1 MaskControl Inference
1:Masked Motion Model ($MMM$), Logits Regularizer ($LR$), mask scheduling function $\gamma(\cdot)$, spatial control signals $s$ (if any), text prompts $W$ (if any).
[…]
3:for all $t$ from $1$ to $T$ do $\triangleright$ Unmask process in $T$ steps
4:    $\{\mathbf{f}\}\leftarrow LR(X_{\overline{\mathbf{M}}},W,s;\phi)$triangleright Logits Regularizer
5:    $l\leftarrow MMM(X_{\overline{\mathbf{M}}},\mathbf{p},\{\mathbf{f}\};\theta)$triangleright Masked Motion Model
6:    for all $i$ from $1$ to $I_{l}$ do $\triangleright$ Logits Optimization
7:         $l_{i+1}=l_{i}-\eta\nabla_{l_{i}}L_{s}(l_{i},s)$
8:    end for
9:    $X_{\overline{\mathbf{M}}}\leftarrow\gamma(l,t)$triangleright mask out tokens based on logits $l$ at time step $t$
10:end for
11:for all $i$ from $1$ to $I_{e}$ do $\triangleright$ Logits Optimization
12:    $e_{c}^{i+1}=e_{c}^{i}-\eta\nabla_{e_{c}^{i}}L_{s}(e_{c}^{i},s)$
13:end for
14:return $Decoder(e_{c})$
A.4 Implementation Details
We modified the MoMask [16] model by retraining it with a cross-entropy loss applied to all tokens, instead of just the masked positions. This retrained model serves as our pretrained base model, and we kept the default hyperparameter settings unchanged. To improve robustness to text variation, we randomly drop 10% of the text conditioning, which also allows the model to be used for Classifier-Free Guidance (CFG). The weight for Eq. 5 is set to $\alpha=0.1$. We use a codebook of size 512, with embeddings of size 512 and 6 residual layers. The Transformer embedding size is set to 384, with 6 attention heads, each with an embedding size of 64, distributed across 8 layers. This configuration demonstrates the feasibility of converting between two different embedding sizes and spaces using the Differentiable Expectation Sampling (DES). The encoder and decoder downsample the motion sequence length by a factor of 4 when mapping to token space. The learning rate follows a linear warm-up schedule, reaching 2e-4 after 2000 iterations, using AdamW optimization. The mini-batch size is set to 512 for training RVQ-VAE and 64 for training the Transformers. During inference, the CFG scale is set to $cfg=4$ for the base layer and $cfg=5$ for the 6 layers of residual, with 10 steps for generation. We use pretrained CLIP model [41] to generate text embeddings, which have a size of 512. These embeddings are then projected down to a size of 384 to match the token size used by the Transformer. Logits Regularizer is a trainable copy of Masked Transformer with the zero linear layer connect to the output each layer of the Masked Transformer. During inference, Logits Optimization applies L2 loss with a learning rate of $0.06$ for 100 iterations in Logits Optimization for each of the 10 generation steps and 600 iterations in the last unmasking step. We apply temperature of 1 for all 10 steps and 1e-8 for residual layers. We follow the implementation from [21, 55, 51], applying the spatial control signal only to joint positions and omitting rotations.
A.5 Full Evaluation on All Joint
Following the evaluation from OmniControl [55], Table 7 extends Table 2 by showing the evaluation for each joint individually. Our MaskControl outperforms SOTA across all metrics. ‘Cross’ is the random combination of joints can be found in Sec. A.13
Table 7: Comparison of text-condition motion generation with spacial control signal on the HumanML3D. The first section, “Train on Pelvis Only,” evaluates our model that was trained solely on the pelvis. The last section, “Train on All Joints”, is trained on all joints and assessing performance for each one. The cross-section reports performance across various combinations of joints.